OpenStudy (anonymous):

the length of a rectangular frame is 5cm more than the width. the area inside the frame is 66 square cm. find the width of the frame I just dnt remember the formula to solve this

OpenStudy (anonymous):

Area = Length * Width Length = W+5 66 = L * W 66=(W+5)W 66 = W^2 +5W solve for W W = 6, L = 11 check: A = L*W 66 = 11*6 66=66 and it's good!

OpenStudy (anonymous):